# SAFE-SQL: Self-Augmented In-Context Learning with Fine-grained Example Selection for Text-to-SQL

### v1

@@ 4.3 Evaluation Metrics
4.4 Performance among SQL difficulty level
We analyze the performance of SAFE-SQL across different SQL difficulty levels and compare it with zero-shot, few-shot prompting methods, and supervised fine-tuning approaches. The results, presented in Table 1, demonstrate that SAFE-SQL achieves overall superior performance, with particularly strong improvements in hard and extra hard categories. Few-shot methods exhibit higher accuracy in Easy and Medium categories, which can be attributed to skeleton-masked selection which retrieves answers directly from the training set, leading to an inflated performance in simpler queries. SAFE-SQL excels in hard and extra hard categories, achieving significantly higher EX. This improvement is notably influenced by the inclusion of reasoning paths, which provide explicit guidance in SQL generation and enhance the model’s ability to construct complex queries, as well as the filtering of misleading examples, which reduces potential confusion and prevents error propagation. While multiple factors contribute to SAFE-SQL’s effectiveness, these mechanisms play a crucial role in enabling the model to generate more accurate and structurally sound SQL queries, especially in challenging scenarios where other approaches struggle.
@@ 4.5 Ablation study
To assess the contribution of each key component in our model, we conduct an ablation study by systematically removing four critical modules: Reasoning Path, Relevance Score, Schema Linking, and Similar Examples. We evaluate the resulting impact on performance using EX shown in Table 2. Our findings indicate that each component plays a crucial role in the model’s effectiveness. Removing the Reasoning Path leads to a 3.5-point drop in EX, highlighting its importance in guiding the model toward generating accurate SQL queries. The absence of the Relevance Score resulted in a 5.8-point decrease in EX, underscoring its contribution to overall performance. Eliminating Schema Linking causes a 7.5-point drop in EX, which demonstrates its critical role in similar example construction. Overall, each of the four components—Reasoning Path, Relevance Score, Schema Linking, and Similar Examples—is essential for achieving optimal performance in SQL generation.
@@ 4.6 Analysis
Effect of Relevance Scoring Thresholds on Performance.
To further evaluate the effectiveness of SAFE-SQL, we conduct a detailed case study using varying thresholds for the relevance scoring mechanism as shown in Figure 4. The self-generated examples are filtered based on relevance scores, with thresholds ranging from 0 to 10. For each test question, the number of high-scoring examples varied due to the specific content and schema structure (e.g., some test questions had six examples with scores $\geq 8$, while others had three). The selected examples are then used during the final inference stage to generate SQL queries. The $\geq 8$ threshold consistently produced the best results, validating the robustness of SAFE-SQL’s relevance score filtering. The results demonstrate that selecting high-quality examples plays a critical role in guiding LLMs to generate accurate SQL queries, regardless of the underlying model.
SQL Question GOLD SQL Query Augmented SQL Question Generated Reasoning Path Relevance Score
@@
Table 4: Examples of original and augmented SQL questions with reasoning paths by GPT-4o.
@@
To assess the impact of the three measuring components—semantic similarity ($\alpha$), keyword & structural similarity ($\beta$), and reasoning path quality ($\gamma$)—on EX, we conduct experiments by varying their respective weightings. The results, presented in Table 5, highlight distinct performance trends across different difficulty levels. Notably, the exclusion of reasoning path quality leads to a drop in EX, particularly in the Hard and Extra Hard. This suggests that a well-structured reasoning path is crucial for handling complex queries, as it provides essential logical steps that bridge the gap between natural language understanding and SQL formulation. Conversely, semantic similarity and structural SQL query similarity have a greater influence on the Easy and Medium levels. This is because these queries tend to be relatively straightforward, meaning that having structurally similar SQL questions in the example set often provides sufficient guidance for generating correct queries. In these cases, direct pattern matching and schema alignment play a larger role. Overall, the findings demonstrate that a balanced combination of all three components is essential for optimizing performance across different levels of query complexity.
@@ 5 Conclusion
In this paper, we introduce SAFE-SQL, a novel unsupervised framework designed to improve Text-to-SQL execution accuracy by generating and filtering self-augmented examples. Through extensive experiments, we show that fine-grained example generation, and optimal threshold filtering contribute to the overall performance increase. Specifically, our method achieves state-of-the-art results, with notable improvements over ablated versions, highlighting the importance of these modules in generating accurate and semantically valid SQL queries. Our findings underscore the capability of large language models when carefully structured and enhanced, to address complex Text-to-SQL tasks.
@@
This research introduces SAFE-SQL, a self-augmented in-context learning framework for Text-to-SQL tasks. While our approach enhances SQL generation without additional fine-tuning, it relies on LLMs, which may inherit biases from training data. We mitigate potential biases and inaccuracies through structured filtering and relevance scoring. Our study uses publicly available datasets, ensuring compliance with data privacy standards. We encourage responsible use of our method, particularly in applications requiring high accuracy and fairness.
@@ Acknowledgement
This research was supported by Institute for Information & Communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program (Chung-Ang University)).
@@ Appendix B Prompts for SAFE-SQL
Semantic Similarity of Questions The selected examples closely match the intent of the test question. Variations in wording do not change the meaning.
@@
Using the selected examples, generate the final SQL query that correctly retrieves the desired result for the given test question. Follow the reasoning patterns observed in the examples. Maintain correct table joins, filters, aggregations, and conditions based on schema constraints. Ensure that the final query accurately represents the intent of the test question while leveraging the insights from the selected examples. Now, generate the final SQL query for the given test question:
@@ Appendix C Impact of model size
Methods Model Type Easy Medium Hard Extra All
SAFE-SQL Llama3.1-8B-Instruct ICL 73.2% 76.1% 63.2% 59.4% 70.5%
SAFE-SQL Deepseek-coder-6.7B ICL 88.8% 65.5% 63.8% 25.3% 64.2%
SAFE-SQL Qwen2.5-7B-Instruct ICL 83.6% 80.7% 78.7% 69.4% 79.2%
SAFE-SQL Starcoder-7B ICL 89.2% 88.9% 84.5% 70.6% 85.2%
Table 10: Execution accuracy performance of different methods across difficulty levels of spider dev set.
D.1 Additional model performance
To evaluate the impact of example generation quality on Text-to-SQL performance, we conducted experiments using different models for final inference. Examples generated by GPT-4o, followed by inference using the target model. Large language models’, such as Qwen 2.5-7B and Deepseek-coder-6.7B, ability to generate high-quality, semantically relevant in-context examples is limited. To mitigate this, we first used GPT-4o to generate in-context examples and filtering examples, then performed final inference using the selected model. Our results show that leveraging GPT-4o for example generation and scoring improved overall execution accuracy by 6.9 points compared to fully relying on Qwen 2.5-7B for the entire process as shown in Table 9. This confirms that high-quality, well-aligned in-context examples play a crucial role in enhancing Text-to-SQL performance, especially in complex queries.

### v2

@@ 4.3 Evaluation Metrics
Method Model Execution Accuracy
Supervised Fine-Tuning (SFT)
Syn-SQL Sense13B 63.4
SQL-Palm Palm 53.6
Zero-shot Methods
Baseline GPT-4 49.2
Baseline GPT-4o 51.8
Baseline GPT-4o-mini 51.2
C3-SQL GPT-4 53.8
Few-shot Methods
Din-SQL GPT-4 55.9
Dail-SQL GPT-4 55.4
ACT-SQL GPT-4 52.8
PTD-SQL GPT-4 57.0
DEA-SQL GPT-4 52.4
Self-augmented In-Context Learning
SAFE-SQL GPT-4 58.9
SAFE-SQL GPT-4o 63.5
SAFE-SQL GPT-4o-mini 62.1
SAFE-SQL Llama3.1-70B-Instruct 60.9
SAFE-SQL Llama3.3-70B-Instruct 61.2
SAFE-SQL Qwen2.5-72B 56.2
SAFE-SQL Gemma3-12B 60.8
SAFE-SQL Gemma3-27B 61.5
Table 2: Execution accuracy on Bird dataset.
4.4 Performance Comparison
Spider Dataset
We analyze the performance of SAFE-SQL across different SQL difficulty levels and compare it with zero-shot, few-shot prompting methods, and supervised fine-tuning approaches. The results, presented in Table 1, demonstrate that SAFE-SQL achieves overall superior performance, with particularly strong improvements in hard and extra hard categories. Few-shot methods exhibit higher accuracy in Easy and Medium categories, which can be attributed to skeleton-masked selection which retrieves answers directly from the training set, leading to an inflated performance in simpler queries. SAFE-SQL excels in hard and extra hard categories, achieving significantly higher EX. This improvement is notably influenced by the inclusion of reasoning paths, which provide explicit guidance in SQL generation and enhance the model’s ability to construct complex queries, as well as the filtering of misleading examples, which reduces potential confusion and prevents error propagation. These multiple factors play a crucial role in enabling the model to generate more accurate and structurally sound SQL queries, especially in challenging scenarios where other approaches struggle. Notably, SAFE-SQL using open-source models such as Gemma3-27B outperforms high-cost methods based on GPT-4, highlighting its cost-effectiveness and strong capability.
Bird Dataset
We also conduct experiments on the Bird Dev dataset in addition to the Spider dataset. Similar to Spider, SAFE-SQL consistently outperforms zero-shot and few-shot methods, achieving 63.5% execution accuracy with GPT-4o, which is even higher than Syn-SQL (63.4%), a supervised fine-tuning approach. This highlights the effectiveness of our SAFE-SQL using a self-augmented in-context learning method.
@@ 4.5 Ablation Study
To assess the contribution of each key component in our model, we conduct an ablation study by systematically removing four critical modules: Reasoning Path, Relevance Score, Schema Linking, and Similar Examples. We evaluate the resulting impact on performance using EX shown in Table 3. Our findings indicate that each component plays a crucial role in the model’s effectiveness. Removing the Reasoning Path leads to a 3.5-point drop in EX, highlighting its importance in guiding the model toward generating accurate SQL queries. The absence of the Relevance Score resulted in a 5.8-point decrease in EX, underscoring its contribution to overall performance. Eliminating Schema Linking causes a 7.5-point drop in EX, which demonstrates its critical role in similar example construction. Overall, each of the four components—Reasoning Path, Relevance Score, Schema Linking, and Similar Examples—is essential for achieving optimal performance in SQL generation, empirically validating our architectural and design choices.
@@ 4.6 Analysis
Inference Time per Query on Spider Dev Set
As shown in Table 1, we compare the inference time of SAFE-SQL with other methods. While baseline methods achieve faster inference via simple zero-shot prompts, they show lower accuracy. Few-shot methods are faster than SAFE-SQL but still underperform in execution accuracy. In contrast, SAFE-SQL leverages example augmentation and filtering process, achieving higher performance with a modest increase in inference time. Despite requiring three LLM calls, SAFE-SQL demonstrates strong zero-shot capabilities without relying on a training set, making the trade-off in latency worthwhile.
@@
GOLD Question GOLD SQL Query Generated Question Reasoning Path Relevance Score
@@
Table 5: Comparative examples of gold-questions and GPT-4o-generated SQL questions, including detailed reasoning paths and the process of computing relevance scores.
@@
Effect of Relevance Scoring Thresholds on Performance.
To further evaluate the effectiveness of SAFE-SQL, we conduct a detailed case study using varying thresholds for the relevance scoring mechanism as shown in Figure 4. The self-generated examples are filtered based on relevance scores, with thresholds ranging from 0 to 10. For each test question, the number of high-scoring examples varied due to the specific content and schema structure (e.g., some test questions have six examples with scores $\geq 8$, while others have three). The selected examples are then used during the final inference stage to generate SQL queries. The $\geq 8$ threshold consistently produces the best results, validating the robustness of SAFE-SQL’s relevance score filtering. The results demonstrate that selecting high-quality examples plays a critical role in guiding LLMs to generate accurate SQL queries, regardless of the underlying model.
@@
To assess the impact of the three measuring components—semantic similarity ($\alpha$), structural similarity ($\beta$), and reasoning path quality ($\gamma$)—on EX, we conduct experiments by varying their respective weightings. The results, presented in Table 6, highlight distinct performance trends across different difficulty levels. Notably, the exclusion of reasoning path quality leads to a drop in EX, particularly in the Hard and Extra Hard. This suggests that a well-structured reasoning path is crucial for handling complex queries, as it provides essential logical steps that bridge the gap between natural language understanding and SQL formulation. Conversely, semantic similarity and structural similarity have a greater influence on the Easy and Medium levels. This is because these queries tend to be relatively straightforward, meaning that having structurally similar SQL questions in the example set often provides sufficient guidance for generating correct queries. In these cases, direct pattern matching and schema alignment play a larger role. Overall, the findings demonstrate that a balanced combination of all three components is essential for optimizing performance across different levels of query complexity.
@@ 5 Conclusion
We introduce SAFE-SQL, a novel unsupervised framework designed for Text-to-SQL. SAFE-SQL generates and filters high-quality self-augmented examples for in-context learning. Extensive experiments demonstrated that both the fine-grained example generation process and optimal threshold filtering contribute significantly to performance gains. Our method achieves state-of-the-art results, showing notable improvements over ablated versions and excelling particularly in challenging extra hard and unseen scenarios.
@@
While our approach enhances SQL generation without additional fine-tuning, it relies on LLMs, which may inherit biases from training data. We mitigate potential biases and inaccuracies through structured filtering and relevance scoring. Our study uses publicly available datasets, ensuring compliance with data privacy standards. We encourage responsible use of our method, particularly in applications requiring high accuracy and fairness.
@@ Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligent Graduate School Program (Chung-Ang University)].
@@ Appendix B Prompts for SAFE-SQL
Semantic Similarity The selected examples closely match the intent of the test question. Variations in wording do not change the meaning.
@@
Using the selected examples, generate the final SQL query that correctly retrieves the desired result for the given test question. Follow the reasoning patterns observed in the examples. Now, generate the final SQL query for the given test question: